# AI-Face: A Million-Scale Demographically Annotated AI-Generated Face Dataset and Fairness Benchmark

### v1

@@ Appendix C Datasheet for AI-Face
C.1 Motivation For Dataset Creation
@@
  • Who funded dataset creation? This work is supported by the U.S. National Science Foundation (NSF) under grant IIS-2348419 and the National Artificial Intelligence Research Resource (NAIRR) Pilot and TACC Lonestar6. Please see Acknowledgment Acknowledgment.
C.2 Data Composition
@@
  • How many instances are there? We include more than 2 million face images from public datasets. Please see Table 13 for details.
@@
  • Is there a label or target associated with each instance? Each image is associated with uncertainty score for gender prediction, uncertainty score for age prediction, uncertainty score for race prediction, gender annotation, age annotation, race annotation, and target label (fake or real).
@@
  • Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description. Yes. Despite our extensive efforts to reduce demographic label noise, including human corrections based on uncertainty scores, there may still be mislabeled instances. Given the dataset’s size of over 2 million images, it is impractical for humans to manually check and correct each image individually.
@@ C.3 Collection Process
  • Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created. The data was collected from February 2024 to April 2024, even though the data were originally released before this time. Please refer to the cited papers in Table 13 for specific original data released time.
C.4 Data Processing
  • Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? Yes. We discussed in ‘Demographically Annotation Generation’ in Section 3.2.
  • Was the ‘raw’ data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the ‘raw’ data. The ‘raw’ data can be acquired through the original data publisher. Please see the cited papers in Table 13.
  • Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point. Yes. We use RetinaFace [60] for detecting and cropping faces to ensure each image only contains one face. Demographic annotations are given by our annotator, see ‘Annotator Development’ in Section 3.1. Our annotator code is available on Our GitHub repository.
@@ C.5 Dataset Distribution
  • When will the dataset be released/first distributed? What license (if any) is it distributed under? The data has been released, under the permissible CC BY-NC-ND 4.0 license for research-based use only. Users can access our dataset by submitting an EULA. Dataset license and EULA is on our GitHub https://github.com/Purdue-M2/AI-Face-FairnessBench.
@@ C.6 Dataset Maintenance
  • Is there a repository to link to any/all papers/systems that use this dataset? Not right now, but we encourage anyone who uses the dataset to cite our paper so it can be easily found. Our fairness benchmark uses this dataset, the code of fairness benchmark is on our GitHub https://github.com/Purdue-M2/AI-Face-FairnessBench.
@@
C.7 Legal and Ethical Considerations
@@
  • Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset? Yes. It is a face image dataset. The age, gender, and race can be identified through the face image, also through the demographic annotation we provide. All of the images that we use are from publicly available data.
C.8 Author Statement and Confirmation of Data License

### v3

@@ Appendix E Datasheet for AI-Face
E.1 Motivation For Dataset Creation
@@
E.2 Data Composition
@@
  • How many instances are there? We include 1,646,545 face images from public datasets. Please see Table B.1 for details.
@@
  • Is there a label or target associated with each instance? Each image is associated with gender annotation, age annotation, skin tone annotation, intersectional attribute (gender and skin tone) annotation, and target label (fake or real).
@@
  • Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description. Yes. Despite our extensive efforts to mitigate the bias that may introduced by the automated annotator and reduce demographic label noise, there may still be mislabeled instances. Given the dataset’s size of over 1 million images and most are generated face images, it is impractical for humans to manually check and correct each image individually.
@@ E.3 Collection Process
  • Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created. The data was collected from February 2024 to April 2024, even though the data were originally released before this time. Please refer to the cited papers in Table B.1 for specific original data released time.
E.4 Data Processing
  • Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? Yes. We discussed in ‘Data Collection’ in Section 3.1.
  • Was the ‘raw’ data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the ‘raw’ data. The ‘raw’ data can be acquired through the original data publisher. Please see the cited papers in Table B.1.
  • Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point. Yes. We use RetinaFace [66] for detecting and cropping faces to ensure each image only contains one face. Demographic annotations are given by our annotator, see ‘Annotation Generation’ in Section 3.2. Our annotator code will not be released considering the ethical guidelines.
@@ E.5 Dataset Distribution
  • When will the dataset be released/first distributed? What license (if any) is it distributed under? The dataset will be released following the paper’s acceptance, and it will be under the permissible CC BY-NC-ND 4.0 license for research-based use only. Users can access our dataset by submitting an EULA.
@@ E.6 Dataset Maintenance
  • Is there a repository to link to any/all papers/systems that use this dataset? Our fairness benchmark uses this dataset, a brief instruction of how to use this dataset and the code of fairness benchmark is on https://github.com/Purdue-M2/AI-Face-FairnessBench.
@@
E.7 Legal and Ethical Considerations
@@
  • Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset? Yes. It is a face image dataset. The age, gender, and skin tone can be identified through the face image, also through the demographic annotation we provide. All of the images that we use are from publicly available data.
E.8 Author Statement and Confirmation of Data License